how many 4-inches by 8 inch bricks are needed to build a walk 6ft. wide and 24ft long.
Answer by Alan3354(30982)

how many 4-inches by 8 inch bricks are needed to build a walk 6ft. wide and 24ft long.
----------
The area of the walk is 6*12 inches by 24*12 inches
Area = 72*288 = 20736 sq inches
Each brick is 4 by 8 = 32 sq inches
20736/32 = 648 bricks